OpenStudy (anonymous):

A number n is multiplied by 5/8 . The product is subtracted from 2/3 . The result is 7/12 . What is n? a. 2/15 b. 6/7 c. 5/12 d. 9/17

OpenStudy (anonymous):

a number \(n\) is multiplied by \(\frac{5}{8}\) which will give \(\frac{5}{8}n\)

OpenStudy (anonymous):

it is then subtracted from \(\frac{2}{3}\) giving \[\frac{2}{3}-\frac{5}{8}n\]

OpenStudy (anonymous):

the answer is given to be \(\frac{7}{12}\) telling you that \[\frac{2}{3}-\frac{5}{8}n=\frac{7}{12}\] and your job is to solve for \(n\) do you know how to do it?

OpenStudy (anonymous):

Not really. Lol

OpenStudy (anonymous):

Can you explain?

OpenStudy (anonymous):

well, if you dislike fractions like i do, then the best thing to do is multiply both sides by the least common multiple of 3, 8 and 12, which is 24 what this really amounts to is cancelling common factors

OpenStudy (anonymous):

if you do not dislike fractions you can work with them and solve in essentially two steps what is your preference?

OpenStudy (anonymous):

I don't like fractions

OpenStudy (anonymous):

ok then lets multiply both sides by 24 first to get rid of them start with \[24\left(\frac{2}{3}-\frac{5}{8}n\right)=\frac{7}{12}\]

OpenStudy (anonymous):

i meant start with \[24\left(\frac{2}{3}-\frac{5}{8}n\right)=24\times \frac{7}{12}\]

OpenStudy (anonymous):

on the right hand side, 12 goes in to 24 two times, so you get \(14\)\[24\left(\frac{2}{3}-\frac{5}{8}n\right)=14\]

OpenStudy (anonymous):

no fraction on the right

OpenStudy (anonymous):

on the left hand side, 3 goes in to 24 eight times, and eight times 2 is 16 also 8 goes in to 24 three times, and three times 5 is 15 this give you \[16-15n=14\]

OpenStudy (anonymous):

subtract 16 from both sides?

OpenStudy (anonymous):

now you can solve in two easy steps subtract 16 from both sides right

OpenStudy (anonymous):

so. youd get -15n= -2

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

final step: divide both sides by \(-15\)

OpenStudy (anonymous):

=7.5

OpenStudy (anonymous):

or 75.

OpenStudy (anonymous):

no, you divided upside down

OpenStudy (anonymous):

you have \(-15n=-2\) divide by \(-15\)

OpenStudy (anonymous):

decimal?

OpenStudy (anonymous):

you should get \[n=\frac{2}{15}\] not \(n=\frac{15}{2}\)

OpenStudy (anonymous):

well, i think given your answer choices you should eschew decimals and just write \(\frac{2}{15}\)

OpenStudy (anonymous):

okay. thanks

OpenStudy (anonymous):

yw